OpenStudy (anonymous):

cos10 cos80 - sin10 sin80

OpenStudy (anonymous):

use "addition angle formula" in reverse \[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\] with \(a=10,b=80\)

OpenStudy (anonymous):

\[\cos90 = 0\]